OpenStudy (anonymous):

Functions f(x) and g(x) are shown below: f(x) = x2 g(x) = x2 + 20x + 100 In which direction and by how many units should f(x) be shifted to obtain g(x)?

OpenStudy (anonymous):

a)right by 5 units b)left by 5 units c)right by 10 units d)left by 10 units

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (ranga):

Complete the square for g(x)

OpenStudy (anonymous):

how do I do that?

OpenStudy (ranga):

g(x) = x^2 + 20x + 100 = (x + 10)^2

OpenStudy (anonymous):

is the answer b?

OpenStudy (ranga):

f(x) = x^2 g(x) = (x + 10)^2 Replacing x by (x+a) in a function, where a is positive, moves the function LEFT by a units. Replacing x by (x-a) in a function, where a is positive, moves the function RIGHT by a units.

OpenStudy (anonymous):

so it would be a then?

OpenStudy (ranga):

How?

OpenStudy (ranga):

f(x) = x^2 g(x) = (x + 10)^2 f(x) becomes g(x) if x is replaced by (x+10). If we replace x by (x+a), the function moves LEFT by "a" units. Here we are replacing x by (x+10). So the function moves LEFT by 10 units.

OpenStudy (anonymous):

oooooohhh so its d